OpenStudy (anonymous):

What is the slope of a line of best fit if its equation is y = −8x + 9? A.1/9 B.-1/8 C.-8 D.9

Nnesha (nnesha):

y = mx + b slope intercept form equation where m is slope which is always with x and b is y-intercept

TheSmartOne (thesmartone):

Hint: In an equation of the format of: \(\sf\Large y=mx+b\) \(\bf\Huge m ~is~the~slope\)

TheSmartOne (thesmartone):

Sooo have you gotten anything? @tykeke878

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

@TheSmartOne

Nnesha (nnesha):

again y = mx + b slope intercept form where m is SLOPE which is always with x m = slope what is m in ur equation ?

OpenStudy (anonymous):

9

Nnesha (nnesha):

is it m ??

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha y = mx + b slope intercept form equation where m is slope which is always with x and b is y-intercept \(\color{blue}{\text{End of Quote}}\) read this again

OpenStudy (anonymous):

8x

Nnesha (nnesha):

not the x x is variable not the slope